OpenStudy (anonymous):

Sqrt52 is between which two whole numbers?

OpenStudy (aum):

7 x 7 = 7^2 = 49 8 x 8 = 8^2 = 64 Since 49 < 52 < 64 7^2 < 52 < 8^2 Take square root throughout: 7 < sqrt(52) < 8

OpenStudy (anonymous):

8 and 7

OpenStudy (anonymous):

Thanks aum showing me the working out to get this answer really helped